The number of married adults in the United States is declining. In 1960, 72% of adults were married. This number had decreased to 51% by 2012. Find the average rate of change in the percent of adults who are married for the years from 1960 to 2012.

Type your detailed solution and:

  • A written description of the steps you use to solve the problem.
  • A mathematical solution to the problem. You should use the Equation Editor to show your math.
Answer #1

We know that the average change is given by the total change divided by the total duration of the time for which the change is considered.

Here, the total change from 1960 to 2012 is 2012 - 1960 = 52 year

Total change in the percentage is given by 51 - 72 = -21%

The minus sign indicates that the percentage has gone down.

Therefore, the average change is given by

-52/21 = -2.476%

It means that the percentage decreases by 2.476 % on an average every year.